# HyPA-RAG: A Hybrid Parameter Adaptive Retrieval-Augmented Generation System for AI Legal and Policy Applications

### v1

@@ Abstract
While Large Language Models (LLMs) excel in text generation and question-answering, their effectiveness in AI legal and policy is limited by outdated knowledge, hallucinations, and inadequate reasoning in complex contexts. Retrieval-Augmented Generation (RAG) systems improve response accuracy by integrating external knowledge but struggle with retrieval errors, poor context integration, and high costs, particularly in interpreting qualitative and quantitative AI legal texts. This paper introduces a Hybrid Parameter-Adaptive RAG (HyPA-RAG) system tailored for AI legal and policy, exemplified by NYC Local Law 144 (LL144). HyPA-RAG uses a query complexity classifier for adaptive parameter tuning, a hybrid retrieval strategy combining dense, sparse, and knowledge graph methods, and an evaluation framework with specific question types and metrics. By dynamically adjusting parameters, HyPA-RAG significantly improves retrieval accuracy and response fidelity. Testing on LL144 shows enhanced correctness, faithfulness, and contextual precision, addressing the need for adaptable NLP systems in complex, high-stakes AI legal and policy applications.  [^1] The demo (Preview in Appendix A.1), dataset and code will be made publicly available upon acceptance of this paper.
@@
Rishi Kalra2, Zekun Wu1,2, Ayesha Gulley1, Airlie Hilliard1, Xin Guan1, Adriano Koshiyama1, Philip Treleaven2 1Holistic AI, 2University College London Correspondence: rishi.kalra.20@ucl.ac.uk, zekun.wu@holisticai.com
@@ 1 Introduction
Recently, the development of Large Language Models (LLMs) capable of processing and generating human-like text has made significant strides in recent years, such as OpenAI’s GPT models (Brown et al., 2020; OpenAI, 2023), Google’s Gemini models (Team et al., 2023) and open alternatives such as the LlaMa series (Touvron et al., 2023a, b; Meta, 2024). These models, which store vast amounts of information within their parameters through extensive pre-training, have demonstrated impressive performance in various tasks, including text generation and question-answering across multiple domains (Brown et al., 2020; Singhal et al., 2023; Wu et al., 2023). Despite this, LLMs encounter limitations when applied to specialised fields such as legal and policy. These include the rapid obsolescence of their knowledge, which is confined to the data available up to the last pre-training date (Yang et al., 2023) and hallucinations, where the model produces text that seems plausible but is factually incorrect or misleading, driven by internal logic rather than actual context (Ji et al., 2022; Huang et al., 2023). Empirical studies show that many AI legal tools overstate their ability to prevent hallucinations (Magesh et al., 2024). Instances of lawyers being penalised for using hallucinated outputs in court documents (Fortune, 2023; Business Insider, 2023) underscore the need for reliable AI question-answering systems in law and policy
Naturally, Retrieval-Augmented Generation (RAG), which enhances LLMs by incorporating external knowledge, is proposed as a solution. However, this comes with its own challenges. Common failure points (Barnett et al., 2024) include missing content, where relevant documents are not retrieved, leading to unanswered questions; context limitations, where retrieved documents are not effectively incorporated into the response generation process due to limitations in consolidation strategies; and extraction failures, where models fail to extract accurate information from the provided context due to noise or conflicting data. Furthermore, advanced retrieval and generation techniques, such as query rewriters and LLM-based quality checkers often result in increased token usage and costs.
@@
To address these challenges, this research integrates three key components (see Figure 6 in Appendix A.2 for a flow overview and Figure 1 for the system design):
Adaptive parameter selection using a domain-specific query complexity classifier to minimise unnecessary token usage,
A hybrid search system combining dense, sparse, and knowledge graph retrieval methods to enhance retrieval accuracy.
An end-to-end evaluation framework that includes the development of a ’gold standard’ dataset, custom question types, and RAG-specific evaluation metrics for robust testing.
These elements are combined to create a hybrid parameter-adaptive RAG system tailored specifically to mitigate the common RAG failure points, for the AI policy domain, using NYC Local Law 144 as the primary corpus. We also provide a streamlit demo for testing purposes.
@@ 2 Background and Related Work
Recent LLM advancements have impacted fields like law and policy, where language complexity and large text volumes are prevalent (Blair-Stanek et al., 2023; Choi et al., 2023; Hargreaves, 2023). LLMs have been used for legal judgment prediction, document drafting, and contract analysis, showing their potential to improve efficiency and accuracy (Shui et al., 2023; Sun, 2023; Šavelka and Ashley, 2023). Techniques like fine-tuning, retrieval augmentation, prompt engineering, and agentic methods have adapted these models for specific legal tasks, enhancing performance in summarisation, drafting, and interpretation (Trautmann et al., 2022; Cui et al., 2023).
Retrieval-Augmented Generation (RAG), as formalized by Lewis et al., enhances pre-trained seq2seq models by integrating external knowledge through indexing, retrieval, and generation stages, improving response specificity and accuracy (Lewis et al., 2020; Gao et al., 2023). RAG systems complement LLMs by combining sparse (e.g., BM25) and dense (e.g., vector) retrieval techniques, using neural embeddings to refine document retrieval and produce grounded, high-quality responses (Jones, 2021; Robertson and Zaragoza, 2009; Devlin et al., 2019; Liu et al., 2019).
To address the limitations of naive RAG, such as insufficient context and retrieval inaccuracies, advanced techniques have been developed, including hybrid retrieval methods, query rewriters, and rerankers to refine relevance (Muennighoff et al., 2022; Ding et al., 2024; Xiao et al., 2023). Hybrid retrieval combines BM25 with semantic embeddings to balance keyword matching and contextual understanding, improving outcomes (Luo et al., 2023; Ram et al., 2022; Arivazhagan et al., 2023). Additionally, knowledge graph retrieval and composed retrievers enhance accuracy and comprehensiveness in document retrieval (Rackauckas, 2024; Sanmartin, 2024; Edge et al., 2024).
Recently, RAG systems have advanced from basic retrieval to dynamic methods involving multi-source integration and domain adaptation (Gao et al., 2023; Ji et al., 2022). Innovations like Self-RAG and KG-RAG improve response quality and minimize hallucinations through adaptive retrieval and knowledge graphs (Asai et al., 2023; Sanmartin, 2024).
Various frameworks have been developed to evaluate RAG systems, including Ragas, which uses reference-free metrics like faithfulness and relevancy (Shahul et al., 2023b). Giskard (AI, 2023) assesses performance using synthetic QA datasets, while ARES utilizes prediction-powered inference (PPI) with specialized LLM judges for accurate evaluation (AI, 2023; Saad-Falcon et al., 2023).
@@ 3 System Design
The hybrid parameter-adaptive RAG system, depicted in Figure 1, integrates vector-based text chunks and a knowledge graph of entities and their relationships to enhance retrieval accuracy. The system employs a hybrid retrieval process, combining sparse (BM25) and dense (vector) methods to retrieve an initial top-$k$ set of results. These results are refined using reciprocal rank fusion based on predefined parameter mappings.
Simultaneously, a knowledge graph retriever identifies relevant triplets, with retrieval depth and keyword selection dynamically adjusted according to query complexity. Results from both BM25 and vector methods are fused again to produce a final optimised set of $k$ chunks.
Optional components include a query rewriter, which generates reformulated queries to improve retrieval. The rewritten queries fetch additional chunks, which are de-duplicated and fused to maintain uniqueness. An optional reranker can further refine chunk ranking if needed. The final set of selected chunks and knowledge graph triplets are then processed within the LLM’s context window for more accurate, contextually relevant responses.
This framework is implemented in two variations: without knowledge graph retrieval, known as Parameter-Adaptive (PA) RAG, and with knowledge graph retrieval, termed Hybrid Parameter-Adaptive (HyPA) RAG.
@@ 4 AI Legal and Policy Corpus
Local Law 144 (LL144) of 2021, enacted by the New York City Department of Consumer and Worker Protection, regulates automated employment decision tools (AEDTs). This paper uses a 15-page version of LL144, including the original legislation and additional context from Subchapter 25. As an early AI-specific law, LL144 is included in the training data of foundational models like GPT-4 and GPT-4o, whose understanding of the law is confirmed manually through targeted prompting and serves as baselines in this research.
LL144 presents significant challenges for AI compliance due to its unique combination of qualitative and quantitative requirements. Unlike most AI legal and policy texts, which are predominantly qualitative, LL144 integrates detailed definitions and procedural guidelines with quantitative compliance metrics. This structure complicates interpretation and retrieval, often exceeding the capabilities of traditional LLMs and RAG systems. Furthermore, AI laws and policies are frequently revised, making them impractical for pre-training and fine-tuning and therefore require a robust method for integrating changes.
@@ 5.1 Dataset Generation
Creating a "gold standard" evaluation set usually requires extensive human expertise and time, but LLMs like GPT-3.5-Turbo can efficiently handle such tasks, if sufficiently prompted. For this purpose, Giskard (AI, 2023) provides a library for synthetic data generation, using LLMs to create various question types from text chunks, such as ’simple’, ’complex’, and ’situational’. We introduce additional types and question generators: ’comparative’, ’complex situational’, ’vague’, and ’rule-conclusion’. Comparative questions require multi-context retrieval to compare concepts. ’Complex situational’ questions involve user-specific contexts and follow-ups. Vague questions obscure parts of the query to test interpretation, while rule-conclusion questions, adapted from LegalBench (Guha et al., 2023), require conclusions based on legislative content. Table 4 in Appendix A.3 summarises these types with examples.
These question generators produce a set of questions, which are then deduplicated. Inaccurate or incomplete questions are identified through a human expert review process, using the criteria outlined in Table 5 in Appendix A.5.
@@ 5.3 Correctness Evaluation
The Spearman coefficient (Figure 2) illustrates how our prompt-based LLM-as-a-judge correctness evaluation aligns with human judgment. Prompts 1 and 2 (Appendix A.7) employ different methods: the baseline prompt provides general scoring guidelines, Prompt 1 offers detailed refinements, and Prompt 2 includes one-shot examples and guidance for edge cases.
@@ 7 Query Complexity Classifier
To enable adaptive parameter selection, we developed a domain-specific query complexity classifier that categorises user queries, each corresponding to specific hyper-parameter mappings. Our analysis of top-$k$ selection indicated different optimal top-$k$ values for various question types, as shown in Figure 7 (Appendix A.4).
@@ 7.3 Classifier Results
Tables 1 and 8 in Appendix A.10 summarise the classification results. We compare performance using macro precision, recall and F1 score. The fine-tuned DistilBERT model achieved the highest F1 scores, 0.90 for the 3-class task and 0.92 for the 2-class task, highlighting the benefits of transfer learning and fine-tuning. The SVM (TF-IDF) and Logistic Regression models also performed well, particularly in binary classification, indicating their effectiveness in handling sparse data. Zero-shot classifiers performed lower, likely due to the lack of task-specific fine-tuning.
@@ 8.2 Hybrid Parameter-Adaptive RAG
Method Faithfulness Answer
Relevancy
Absolute
Correctness (1-5)
Correctness
(Threshold=4.0)
$k$ 0.7723 0.7940 4.0409 0.7621
$k$, $Q$ 0.8971 0.7778 4.2528 0.8141
$k$, $Q$ + reranker 0.9098 0.7902 4.2342 0.8178
$k$, $K^{*}$, $S^{*}$ 0.8733 0.7635 4.1227 0.8141
$k$, $K$, $S$ 0.8660 0.7780 4.1822 0.8030
$k$, $K$, $S$ + reranker 0.8821 0.7872 4.1858 0.8178
$k$, $K$, $S$, $Q$ 0.8465 0.7734 4.1338 0.7918
$k$, $K$, $S$, $Q$ + reranker 0.8689 0.7853 4.1859 0.8402
@@ 8.3 RAG Results
The adaptive methods generally outperform the fixed $k$ baseline across most metrics (Table 2). PA-RAG with $k,Q$ (2 class) achieves the highest faithfulness score of 0.9044, which is an improvement of 0.0564 over the best fixed $k=10$ method (0.8480). Similarly, the PA $k,Q$ (3 class) configuration also performs strongly with a faithfulness score of 0.8971, surpassing all fixed $k$ methods.
For answer relevancy, the PA $k,Q$ (2 class) model achieves a score of 0.7910, which is nearly on par with the best fixed $k=10$ method at 0.7917, showing a slight difference of 0.0007. The PA $k,Q$ (3 class) model has a relevancy score of 0.7778, a drop of 0.0139 compared to the best fixed method.
In terms of absolute correctness, both PA models, $k,Q$ (2 class) and $k,Q$ (3 class), achieve scores of 4.2491 and 4.2528, respectively, which are improvements of approximately 0.1896 and 0.1933 over the best fixed method ($k=10$) score of 4.0595. This suggests that adaptive parameter settings significantly enhance the model’s ability to provide correct answers.
Correctness scores also favour the adaptive methods. PA $k,Q$ (3 class) model reaches a score of 0.8141, which is 0.0483 higher than the best fixed $k=10$ score of 0.7658. PA $k,Q$ (2 class) model shows similar strength with a score of 0.8104. HyPA show more varied results. HyPA $k,Q,K,S$ (2 class) achieves a correctness score of 0.7770, a modest increase of 0.0112 over the fixed $k=7$, suggesting room for further optimisation.
@@ 8.4 System Ablation Study
We evaluate the impact of adaptive parameters, a reranker (bge-reranker-large), and a query rewriter on model performance using PA and HyPA RAG methods with 2-class (Table 9 in Appendix A.11) and 3-class classifiers (Table 3).
The highest Answer Relevancy (0.7940) is achieved by varying $k$ alone, suggesting that simpler, focused responses facilitate question reconstruction. The $k$, $Q$ + reranker configuration achieves a slightly lower relevancy score (0.7902), indicating that query rewriting and reranking, while enhancing other metrics, introduce complexity that marginally reduces clarity.
The $k$, $Q$ + reranker configuration also achieves the highest Faithfulness (0.9098), showing that combining adaptive top-$k$ selection with query rewriting and reranking improves factual consistency. This setup provides high Absolute Correctness (4.2342), although slightly lower than $k$, $Q$ alone (4.2528), indicating that while reranking improves response quality, it may slightly decrease overall accuracy. However, the Correctness Score improves from 0.8141 to 0.8178, highlighting an increase in responses classified as "correct" (scores of 4 or higher).
Adding a knowledge graph in the $k$, $K$, $S$ configuration maintains the same Correctness Score (0.8141) as $k$, $Q$ but reduces Absolute Correctness by 0.1301, suggesting added complexity might lower overall answer quality.
While the $k$, $K$, $S$, $Q$ + reranker configuration does not lead in Faithfulness, Answer Relevancy, or Absolute Correctness, it achieves the highest Correctness Score (0.8402), outperforming $k$, $Q$ + reranker by 0.0224, demonstrating the effectiveness of adaptive parameters and reranking in meeting the correctness threshold.

### v2

@@ Abstract
Large Language Models (LLMs) face limitations in AI legal and policy applications due to outdated knowledge, hallucinations, and poor reasoning in complex contexts. Retrieval-Augmented Generation (RAG) systems address these issues by incorporating external knowledge, but suffer from retrieval errors, ineffective context integration, and high operational costs. This paper presents the Hybrid Parameter-Adaptive RAG (HyPA-RAG) system, designed for the AI legal domain, with NYC Local Law 144 (LL144) as the test case. HyPA-RAG integrates a query complexity classifier for adaptive parameter tuning, a hybrid retrieval approach combining dense, sparse, and knowledge graph methods, and a comprehensive evaluation framework with tailored question types and metrics. Testing on LL144 demonstrates that HyPA-RAG enhances retrieval accuracy, response fidelity, and contextual precision, offering a robust and adaptable solution for high-stakes legal and policy applications.
@@
Rishi Kalra1,2, Zekun Wu1,2thanks: Corresponding author, Ayesha Gulley1, Airlie Hilliard1, Xin Guan1, Adriano Koshiyama1, Philip Treleaven211footnotemark: 1 1Holistic AI, 2University College London
@@ 1 Introduction
Large Language Models (LLMs) like GPT (Brown et al., 2020; OpenAI, 2023), Gemini (Team et al., 2023), and Llama (Touvron et al., 2023a, b; Meta, 2024) have advanced question answering across domains (Brown et al., 2020; Singhal et al., 2023; Wu et al., 2023). However, they face challenges in domains like law and policy due to outdated knowledge limited to pre-training data (Yang et al., 2023) and hallucinations, where outputs appear plausible but are factually incorrect (Ji et al., 2022; Huang et al., 2023). Empirical evidence indicates that many AI tools for legal applications overstate their ability to prevent hallucinations (Magesh et al., 2024). Cases of lawyers penalized for using hallucinated court documents (Fortune, 2023; Business Insider, 2023) highlight the need for reliable AI systems in legal and policy contexts.
Retrieval-Augmented Generation (RAG) integrates external knowledge into LLMs to address their limitations but faces challenges. These include missing content, where relevant documents are not retrieved; context limitations, where retrieved documents are poorly integrated into responses; and extraction failures due to noise or conflicting data (Barnett et al., 2024). Advanced techniques like query rewriters and LLM-based quality checks improve quality but increase token usage and costs.
@@
This research presents the Hybrid Parameter-Adaptive RAG (HyPA-RAG) system to address RAG challenges in AI policy, using NYC Local Law 144 as a test corpus. HyPA-RAG includes adaptive parameter selection with a query complexity classifier to reduce token usage, a hybrid retrieval system combining dense, sparse, and knowledge graph methods to improve accuracy, and an evaluation framework with a gold dataset, custom question types, and RAG-specific metrics. These components address common RAG failures and enhance AI applications in legal and policy domains.
@@ 2 Background and Related Work
Recent LLM advancements have influenced law and policy, where complex language and large text volumes are common (Blair-Stanek et al., 2023; Choi et al., 2023; Hargreaves, 2023). LLMs have been applied to legal judgment prediction, document drafting, and contract analysis, improving efficiency and accuracy (Shui et al., 2023; Sun, 2023; Šavelka and Ashley, 2023). Techniques like fine-tuning, retrieval augmentation, prompt engineering, and agentic methods have further enhanced performance in summarization, drafting, and interpretation (Trautmann et al., 2022; Cui et al., 2023).
RAG enhances language models by integrating external knowledge through indexing, retrieval, and generation, using sparse (e.g., BM25) and dense (e.g., vector) techniques with neural embeddings to improve response specificity, accuracy, and grounding (Lewis et al., 2020; Gao et al., 2023; Jones, 2021; Robertson and Zaragoza, 2009; Devlin et al., 2019; Liu et al., 2019). To overcome naive RAG’s limitations, such as poor context and retrieval errors, advanced methods like hybrid retrieval, query rewriters, and rerankers have been developed (Muennighoff et al., 2022; Ding et al., 2024; Xiao et al., 2023). Hybrid retrieval combines BM25 with semantic embeddings for better keyword matching and contextual understanding (Luo et al., 2023; Ram et al., 2022; Arivazhagan et al., 2023), while knowledge graph retrieval and composed retrievers improve accuracy and comprehensiveness (Rackauckas, 2024; Sanmartin, 2024; Edge et al., 2024). Recently, RAG systems have advanced from basic retrieval to dynamic methods involving multi-source integration and domain adaptation (Gao et al., 2023; Ji et al., 2022). Innovations like Self-RAG and KG-RAG improve response quality and minimize hallucinations through adaptive retrieval and knowledge graphs (Asai et al., 2023; Sanmartin, 2024). Frameworks for evaluating RAG systems include Ragas, which uses reference-free metrics like faithfulness and relevancy (Shahul et al., 2023b), Giskard, which leverages synthetic QA datasets (AI, 2023), and ARES, which employs prediction-powered inference with LLM judges for precise evaluation (AI, 2023; Saad-Falcon et al., 2023).
@@ 3 System Design
The Hybrid Parameter-Adaptive RAG (HyPA-RAG) system, shown in Figure 1, integrates vector-based text chunks and a knowledge graph of entities and relationships to improve retrieval accuracy. It employs a hybrid retrieval process that combines sparse (BM25) and dense (vector) methods to retrieve an initial top-$k$ set of results, refined using reciprocal rank fusion based on predefined parameter mappings. A knowledge graph (KG) retriever dynamically adjusts retrieval depth and keyword selection based on query complexity, retrieving relevant triplets. Results are combined with the KG results appending it to the retrieved chunks to generate an final set of $k$ chunks. Optional components include a query rewriter to enhance retrieval with reformulated queries and a reranker for further refining chunk ranking. De-duplicated rewritten query results are integrated into the final set, which, along with knowledge graph triplets, is processed within the LLM’s context window for precise, contextually relevant responses. The framework has two variations: Parameter-Adaptive (PA) RAG, which excludes knowledge graph retrieval, and Hybrid Parameter-Adaptive (HyPA) RAG, which incorporates it.
@@ 4 AI Legal and Policy Corpus
Local Law 144 (LL144) of 2021, enacted by New York City’s Department of Consumer and Worker Protection (DCWP), regulates automated employment decision tools (AEDTs). This study uses a 15-page version of LL144, combining the original law with DCWP enforcement rules. As an early AI-specific law, LL144 is included in GPT-4 and GPT-4o training data, verified via manual prompting, and serves as a baseline in this research. The complexity of LL144 motivates our system’s design for several reasons: (1) it requires multi-step reasoning and concept linking due to its mix of qualitative and quantitative requirements—definitions, procedural guidelines, and compliance metrics—that semantic similarity alone cannot capture, addressed through our knowledge graph; (2) seemingly simple queries can be ambiguous or require multiple information chunks, making a query rewriter and classifier necessary; and (3) while not specific to our adaptive classifier, the evolving nature of AI laws limits the effectiveness of static pre-training, making retrieval-augmented systems better suited to handle frequent updates. These factors go beyond what standard LLMs and basic RAG systems can manage, justifying the need for our approach.
@@ 5.1 Dataset Generation
We created a "gold standard" evaluation set to assess system performance, leveraging GPT-3.5-Turbo and Giskard (AI, 2023) for efficient question generation. The dataset includes various question types, such as ’simple’, ’complex’, ’situational’, and novel types like ’comparative’, ’complex situational’, ’vague’, and ’rule-conclusion’ (inspired by LegalBench (Guha et al., 2023)). These questions test multi-context retrieval, user-specific contexts, query interpretation, and legal reasoning. Generated questions were deduplicated and refined through expert review to ensure accuracy and completeness, using the criteria outlined in Table 4 in Appendix A.5.
@@ 5.3 Correctness Evaluation
The Spearman coefficient (Figure 2) shows how our prompted LLM correctness judge aligns with human judgment. Prompts 1 and 2 (Appendix A.7) employ different methods: the baseline prompt provides general scoring guidelines, Prompt 1 offers detailed refinements, and Prompt 2 includes one-shot examples and edge cases.
@@ 7 Query Complexity Classifier
We developed a domain-specific query complexity classifier for adaptive parameter selection, mapping queries to specific hyper-parameters. Unlike Adaptive RAG (jeong-etal-2024-adaptive), our classifier influences not only the top-$k$ but also knowledge graph and query rewriter parameters. Our analysis of top-$k$ selection indicated different optimal top-$k$ values for various question types, as shown in Figure 6 (Appendix A.4).
@@ 7.3 Classifier Results
Tables 1 and 7 in Appendix A.10 summarise the classification results. We compare performance using macro precision, recall and F1 score. The fine-tuned DistilBERT model achieved the highest F1 scores, 0.90 for the 3-class task and 0.92 for the 2-class task, highlighting the benefits of transfer learning and fine-tuning. The SVM (TF-IDF) and Logistic Regression models also performed well, particularly in binary classification, indicating their effectiveness in handling sparse data. Zero-shot classifiers performed lower.
@@ 8.3 RAG Results
Adaptive methods consistently outperform fixed $k$ baselines. PA-RAG $k,Q$ (2 class) achieves the highest faithfulness score of 0.9044, a 0.0564 improvement over the best fixed method ($k=10$, 0.8480). Similarly, PA $k,Q$ (3 class) achieves 0.8971, surpassing all fixed $k$ methods. For answer relevancy, PA $k,Q$ (2 class) scores 0.7910, nearly matching the best fixed method (0.7917), while PA $k,Q$ (3 class) scores slightly lower at 0.7778. In absolute correctness, PA $k,Q$ (2 class) and $k,Q$ (3 class) achieve 4.2491 and 4.2528, respectively, improving by 0.1896 and 0.1933 over the best fixed method ($k=10$, 4.0595). Correctness scores further highlight the advantage, with PA $k,Q$ (3 class) scoring 0.8141, 0.0483 higher than the fixed baseline (0.7658). HyPA results are more variable. HyPA $k,Q,K,S$ (2 class) achieves a correctness score of 0.7770, a modest 0.0112 improvement over fixed $k=7$, indicating potential for further optimization.
@@ 8.4 System Ablation Study
We evaluate the impact of adaptive parameters, a reranker (bge-reranker-large), and a query rewriter on model performance using PA and HyPA RAG methods with 2-class (Table 9 in Appendix A.12) and 3-class classifiers (Table 8 in Appendix A.11).
Adaptive parameters, query rewriting, and reranking significantly influence RAG performance. Varying the top-$k$ chunks alone achieves the highest Answer Relevancy (0.7940), while adapting the top-$k$ and number of query rewrites with a reranker ($k$, $Q$ + reranker) delivers the highest Faithfulness (0.9098) and improves Correctness Score from 0.8141 to 0.8178. Adding a knowledge graph ($k$, $K$, $S$) maintains the same Correctness Score (0.8141) but lowers Absolute Correctness. The HyPA ($k$, $K$, $S$, $Q$ + reranker) setup achieves the highest Correctness Score (0.8402), showing the value of adaptive parameters and reranking in improving correctness.